# Hyperspectral Pansharpening: Critical Review, Tools and Future Perspectives

### v1

@@ IV-B Is this worth the effort?
As noted above, in the absence of suitable datasets, a common practice for DL-based methods is to use a single image split into two non-overlapping parts for training and testing. To simulate this condition, we train our 5 supervised models, following the original experimental configurations, on a fraction (about 3/4) of the Udine image, listed in Tab.V(B), and test them on the remaining part, call these parts Udine/1 and Udine/2. Numerical results are reported in Tab. VI (only ERGAS for brevity). All five methods work quite well (on visual inspection) with very close ERGAS indicators, ranging from 1.7 to 2.1. Then, we train the same models on the proposed training set, see Tab. V(A), obtaining slightly worse results in all cases [^4] Much worse for DHP-DARN, which seems to be an outlier in this case, i.e., extremely sensitive to the training set.. As expected the statistical “alignment” between training and testing data offers some advantages. In fact, even if Udine/1 and Udine/2 are disjoint, they come from the same acquisition and share the same sensing geometry, daylight and atmosphere conditions, land cover, etc.
Things change completely when models are tested on a new image, Cagliari, again in Tab.V(B). The models trained on the proposed dataset exhibit a rather stable performance, with a slight average improvement concerning Udine/2. Neither Udine nor Cagliari contribute to the training set, so these results suggest that the latter can more easily pansharpened because of its simpler structure (see also Fig. 2). However, the models trained on Udine/1 suffer a catastrophic impairment, with ERGAS increasing by more than 80% on average. This is because their weights, over-fit to a small training set, are unable to generalize well to new data. In summary, this experiment, although limited in scope, confirms that the use of large, well-structured datasets improves generalization and prevents biases in experimental results.
TABLE VI: ERGAS scores with small/aligned vs large training datasets.
@@
TABLE VII: Results at Reduced Resolution. Best, Top-5 and Worst-5 are in bold green, green and red. The average score is in the rightmost column.
@@
TABLE VIII: Results at Full Resolution. Best, Top-5 and Worst-5 are in bold green, green and red. The average score is in the rightmost column.
@@ V Experimental results
Let us now delve into the numerical results summarized in Tab. VII and Tab. VIII for RR and FR datasets, respectively. For easier understanding, we have highlighted the 5 best and 5 worst performing techniques in green and red, respectively, with the best one emphasized in bold. Comments to the results are gathered and ordered according to the above list of quality features.
@@ V-A Generalization across datasets
To assess a method’s ability to generalize across datasets, we fix the scale and index, e.g., ERGAS, reduced resolution, and analyze score variations across images. At reduced resolution (see Tab. VII) virtuous examples are MTF-GLP-FS and MTF-GLP-HPM-R, which exhibit high average scores for all three indexes, consistently across all test images. The same also holds for some DL-based methods, e.g., Hyper-DSNet, R-PNN, and others with worse average scores. Examples of methods with generalization issues at reduced resolution are MTF-GLP-HPM or PCA-Z-PNN. The former gets very good scores on Cagliari followed by bad ones on Udine. The latter performs well on Tabasco, but much worse on Ford.
Moving to the full resolution results in Tab. VIII, consistent behaviour is observed across the four test scenes for almost all methods on the spectral distortion index $D_{\lambda}$. However, while all reduced resolution indices can be considered alternative measures of overall quality, the full resolution indices $D_{\lambda}$ and $D_{S}$ each focus on a specific feature, i.e. spectral or spatial quality. Therefore, they are more likely to be uniform across images, following the characteristics of the method. Inconsistencies are more easily observed in the RQNR hybrid index since it results from the combination of $D_{\lambda}$ and $D_{S}$. This is the case, for example, of BT-H or BDSD-PC, which show opposite spectral and spatial behaviours resulting in unstable RQNR rankings.
@@ V-B Generalization across scales
To simplify the analysis of results, we consider only average scores, focusing, for each method, on the coherence between rankings registered at RR with those occurring at FR. The results of Tab. VII and Tab. VIII fully confirm that the methods generalize much better across different scenes than across scales. Some methods, like PRACS and PCA-Z-PNN, which struggle a bit at RR, seem to work well on FR data, while others, in particular some supervised DL-based method, suffer a performance hit when switching to FR tests. Examples of methods that tend to perform fairly well (but not top) at both scales are GSA, Hyper-DSNet and R-PNN. From these numbers, a picture emerges in which most methods seem to be optimized for just one scale, suffering on the other. In some cases, these results are the effect of an explicit design choice, as in the case of supervised DL methods, where training must be carried out in the RR domain. Given this apparent compromise, that is, working well only at one scale, one might wonder which one is preferable. On the one hand, the ultimate goal is to work well on FR data. On the other hand, at this scale, only consistency measures and not objective quality indicators can be observed. Eventually, the question remains open.
@@ V-C Joint spectral-spatial quality
In fusing the PAN and HS components we aim to keep both the spatial detail of the former and the spectral richness of the latter. However, these goals are somewhat conflicting, and sometimes one property is obtained by sacrificing the other. To analyze this trade-off we can refer to the FR consistency indices, $D_{\lambda}$ and $D_{S}$, which evaluate spatial and spectral quality separately. Indeed, the FR results of Tab. VIII highlight several situations where this undesirable behaviour emerges clearly, with methods that show high scores on $D_{\lambda}$ but poor performance on $D_{S}$, or viceversa. The hybrid RQNR score, which weights the two indicators, can help evaluate the correct balance achieved by a method, with poor balance, signalled by a low RQNR, indicating ineffective fusion. Experimental results show that some recent DL-based methods, like DIP-HyperKite, R-PNN and PCA-Z-PNN, offer excellent RQNR scores. However, quite dated classical methods, like GSA and PRACS, are no less competitive. This confirms the observation already made in the introduction, that DL has yet to realize its full potential for HS pansharpening.
@@ V-D Subjective visual quality
For the RR framework we selected crops from Udine (see Fig. 3) and Ford Country (see Fig. 4), showing results for visible bands on the top and NIR-SWIR bands on the bottom. On the top-left we show the ground truth (namely, the original HS), and then the pansharpening results for all methods. For both scenes, visual inspection seems to confirm the indications provided by the numerical results of Tab. VII. Without going into too much detail, it can be easily verified that images that appear too blurred or present spectral distortion correspond to methods ranked in the worst positions in Tab. VII. On the other hand, given the availability of the GT, numerical assessment is very reliable in this domain.
In the FR framework, visual inspection turns out to be much more informative. Here, lacking the GT, numerical results are less reliable as they only check consistency with the two input components, PAN and HS. Visual results are displayed in Figures 5-8 for crops taken from Cagliari and Tabasco, analyzed in the visible and NIR-SWIR ranges, respectively. In all cases, the original PAN and HS (the latter expanded to match sizes) are shown on the top left, followed by the output of all methods. The comparison with the original HS shows some significant spectral deviations (colours do not match) like for HyperPNN and DHP-DARN on Cagliari. In general, these problems were correctly predicted by the numerical indications of Tab. VIII). In most cases, the spectral quality is definitely good, both in the visible and NIR-SWIR ranges, confirming that spectral fidelity is not the main issue in HS pansharpening. The situation is more critical for spatial fidelity. By comparing results with the PAN, several problems emerge, concerning blurring or, on the opposite side, oversharpening. Some methods (PRACS, HySURE, SR-D, TV) produce systematic blurred outputs on both Cagliari and Tabasco and both in the visible and NIR-SWIR ranges. In some other cases, this is a spectrally selective phenomenon: for example, BDSD-PC blurs in the NIR-SWIR while AWLP blur in the visible range. Finally, in a few cases (HSpeNet, DIP-HyperKite) the blur is image-dependent, occurring only on Cagliari. In the opposite direction, some methods deliver over-sharpened output images, especially BT-H in the NIR-SWIR range.
@@ V-E Computational Time
All experiments were carried out on the same server, an NVIDIA DGX Station A100, equipped with a 64-core AMD EPYC 7742 processor, 504 GB of DDR4 RAM, and four NVIDIA A100-SXM4 GPUs with 40 GB of GDDR5 memory each. The conventional CS, MRA, and MBO methods were run on the CPU, while the DL-based methods were run on a single GPU. The average run time for each method on the 600$\times$600 RR test images is reported in Fig. 9. The run times for the 4 times larger (1200$\times$1200) FR test images scale linearly with the size and hence are not reported. However, computational scalability depends on the available hardware, so we do not draw general conclusions about that. Some methods, both DL-based and not, have negligible run time, a quality to be considered in practical applications. Others are exceedingly slow, such as the CPU-based BDSD-PC and the GPU-based DHP-DARN. In general, all DL-based methods that try to adapt to the variability of input images pay a price in terms of computational complexity, DHP-DARN and DIP-HyperKite to estimate the deep image prior, and R-PNN and PCA-Z-PNN for the target-adaptation phase.

### v2

@@ IV-B Is this worth the effort?
As noted above, in the absence of suitable datasets, a common practice for DL-based methods is to use a single image split into two non-overlapping parts for training and testing. To simulate this condition, we train our 5 supervised models, following the original experimental configurations, on a fraction (about 3/4) of the Udine image, listed in Tab.VI(B), and test them on the remaining part, call these parts Udine/1 and Udine/2. Numerical results are reported in Tab. VII (only ERGAS for brevity). All five methods work quite well (on visual inspection) with very close ERGAS indicators, ranging from 1.7 to 2.1. Then, we train the same models on the proposed training set, see Tab. VI(A), obtaining slightly worse results in all cases [^4] Much worse for DHP-DARN, which seems to be an outlier in this case, i.e., extremely sensitive to the training set.. As expected the statistical “alignment” between training and testing data offers some advantages. In fact, even if Udine/1 and Udine/2 are disjoint, they come from the same acquisition and share the same sensing geometry, daylight and atmosphere conditions, land cover, etc.
Things change completely when models are tested on a new image, Cagliari, again in Tab.VI(B). The models trained on the proposed dataset exhibit a rather stable performance, with a slight average improvement concerning Udine/2. Neither Udine nor Cagliari contribute to the training set, so these results suggest that the latter can more easily pansharpened because of its simpler structure (see also Fig. 2). However, the models trained on Udine/1 suffer a catastrophic impairment, with ERGAS increasing by more than 80% on average. This is because their weights, over-fit to a small training set, are unable to generalize well to new data. In summary, this experiment, although limited in scope, confirms that the use of large, well-structured datasets improves generalization and prevents biases in experimental results.
TABLE VII: ERGAS scores with small/aligned vs large training datasets.
@@
TABLE VIII: Results at Reduced Resolution. Best, Top-5 and Worst-5 are in bold green, green and red. The average score is in the rightmost column.
@@
TABLE IX: Results at Full Resolution. Best, Top-5 and Worst-5 are in bold green, green and red. The average score is in the rightmost column.
@@ V Experimental results
Let us now delve into the numerical results summarized in Tab. VIII and Tab. IX for RR and FR datasets, respectively. For easier understanding, we have highlighted the 5 best and 5 worst performing techniques in green and red, respectively, with the best one emphasized in bold. Comments to the results are gathered and ordered according to the above list of quality features.
@@ V-A Generalization across datasets
To assess a method’s ability to generalize across datasets, we fix the scale and index, e.g., ERGAS, reduced resolution, and analyze score variations across images. At reduced resolution (see Tab. VIII) virtuous examples are MTF-GLP-FS and MTF-GLP-HPM-R, which exhibit high average scores for all three indexes, consistently across all test images. The same also holds for some DL-based methods, e.g., Hyper-DSNet, R-PNN, and others with worse average scores. Examples of methods with generalization issues at reduced resolution are MTF-GLP-HPM or PCA-Z-PNN. The former gets very good scores on Cagliari followed by bad ones on Udine. The latter performs well on Tabasco, but much worse on Ford.
Moving to the full resolution results in Tab. IX, consistent behaviour is observed across the four test scenes for almost all methods on the spectral distortion index $D_{\lambda}$. However, while all reduced resolution indices can be considered alternative measures of overall quality, the full resolution indices $D_{\lambda}$ and $D_{S}$ each focus on a specific feature, i.e. spectral or spatial quality. Therefore, they are more likely to be uniform across images, following the characteristics of the method. Inconsistencies are more easily observed in the RQNR hybrid index since it results from the combination of $D_{\lambda}$ and $D_{S}$. This is the case, for example, of BT-H or BDSD-PC, which show opposite spectral and spatial behaviours resulting in unstable RQNR rankings.
@@ V-B Generalization across scales
To simplify the analysis of results, we consider only average scores, focusing, for each method, on the coherence between rankings registered at RR with those occurring at FR. The results of Tab. VIII and Tab. IX fully confirm that the methods generalize much better across different scenes than across scales. Some methods, like PRACS and PCA-Z-PNN, which struggle a bit at RR, seem to work well on FR data, while others, in particular some supervised DL-based method, suffer a performance hit when switching to FR tests. Examples of methods that tend to perform fairly well (but not top) at both scales are GSA, Hyper-DSNet and R-PNN. From these numbers, a picture emerges in which most methods seem to be optimized for just one scale, suffering on the other. In some cases, these results are the effect of an explicit design choice, as in the case of supervised DL methods, where training must be carried out in the RR domain. Given this apparent compromise, that is, working well only at one scale, one might wonder which one is preferable. On the one hand, the ultimate goal is to work well on FR data. On the other hand, at this scale, only consistency measures and not objective quality indicators can be observed. Eventually, the question remains open.
@@ V-C Joint spectral-spatial quality
In fusing the PAN and HS components we aim to keep both the spatial detail of the former and the spectral richness of the latter. However, these goals are somewhat conflicting, and sometimes one property is obtained by sacrificing the other. To analyze this trade-off we can refer to the FR consistency indices, $D_{\lambda}$ and $D_{S}$, which evaluate spatial and spectral quality separately. Indeed, the FR results of Tab. IX highlight several situations where this undesirable behaviour emerges clearly, with methods that show high scores on $D_{\lambda}$ but poor performance on $D_{S}$, or viceversa. The hybrid RQNR score, which weights the two indicators, can help evaluate the correct balance achieved by a method, with poor balance, signalled by a low RQNR, indicating ineffective fusion. Experimental results show that some recent DL-based methods, like DIP-HyperKite, R-PNN and PCA-Z-PNN, offer excellent RQNR scores. However, quite dated classical methods, like GSA and PRACS, are no less competitive. This confirms the observation already made in the introduction, that DL has yet to realize its full potential for HS pansharpening.
@@ V-D Subjective visual quality
For the RR framework we selected crops from Udine (see Fig. 3) and Ford Country (see Fig. 4), showing results for visible bands on the top and NIR-SWIR bands on the bottom. On the top-left we show the ground truth (namely, the original HS), and then the pansharpening results for all methods. For both scenes, visual inspection seems to confirm the indications provided by the numerical results of Tab. VIII. Without going into too much detail, it can be easily verified that images that appear too blurred or present spectral distortion correspond to methods ranked in the worst positions in Tab. VIII. On the other hand, given the availability of the GT, numerical assessment is very reliable in this domain.
In the FR framework, visual inspection turns out to be much more informative. Here, lacking the GT, numerical results are less reliable as they only check consistency with the two input components, PAN and HS. Visual results are displayed in Figures 5-8 for crops taken from Cagliari and Tabasco, analyzed in the visible and NIR-SWIR ranges, respectively. In all cases, the original PAN and HS (the latter expanded to match sizes) are shown on the top left, followed by the output of all methods. The comparison with the original HS shows some significant spectral deviations (colours do not match) like for HyperPNN and DHP-DARN on Cagliari. In general, these problems were correctly predicted by the numerical indications of Tab. IX). In most cases, the spectral quality is definitely good, both in the visible and NIR-SWIR ranges, confirming that spectral fidelity is not the main issue in HS pansharpening. The situation is more critical for spatial fidelity. By comparing results with the PAN, several problems emerge, concerning blurring or, on the opposite side, oversharpening. Some methods (PRACS, HySURE, SR-D, TV) produce systematic blurred outputs on both Cagliari and Tabasco and both in the visible and NIR-SWIR ranges. In some other cases, this is a spectrally selective phenomenon: for example, BDSD-PC blurs in the NIR-SWIR while AWLP blur in the visible range. Finally, in a few cases (HSpeNet, DIP-HyperKite) the blur is image-dependent, occurring only on Cagliari. In the opposite direction, some methods deliver over-sharpened output images, especially BT-H in the NIR-SWIR range.
@@ V-E Computational Time
All experiments were carried out on the same server, an NVIDIA DGX Station A100, equipped with a 64-core AMD EPYC 7742 processor, 504 GB of DDR4 RAM, and four NVIDIA A100-SXM4 GPUs with 40 GB of GDDR5 memory each. The conventional CS, MRA, and MBO methods were run on the CPU, while the DL-based methods were run on a single GPU. The average run time for each method on the 600$\times$600 RR test images is reported in Fig. 9. The run times for the 4 times larger (1200$\times$1200) FR test images scale linearly with the size and hence are not reported. However, computational scalability depends on the available hardware, so we do not draw general conclusions about that. Some methods, both DL-based and not, have negligible run time, a quality to be considered in practical applications. Others are exceedingly slow, such as the CPU-based BDSD-PC and the GPU-based DHP-DARN. In general, all DL-based methods that try to adapt to the variability of input images pay a price in terms of computational complexity, DHP-DARN and DIP-HyperKite to estimate the DIP, and R-PNN and PCA-Z-PNN for the target-adaptation phase.
V-F To what extent existing MS pansharpening methods generalize to the HS case?
Given the nature of the problem at hand, it is fair to ask to what extent a pansharpening method conceived for the PAN-MS fusion can keep working satisfactory when replacing the MS with a HS datacube. Skipping implementation issues due to the adaptation and computational implications that differ from one method to another, a partial answer on quality can now be given on the basis of the achieved results. The unique challenges present in the HS case are clear: much larger (and variable) number of bands; spectral separation between the PAN and a large (the majority) number of bands; larger resolution ratios to face (e.g., 6 for PRISMA); larger volume of data per geographic area.
Despite these differences, the similarities between the two problems have led many researchers to apply MS pansharpening methods to HS data, either directly or with minor modifications. From the analysis of results, both quantitatively (refer to Table VIII and Table IX) and qualitatively (refer to Figures 5-8), it appears that some MS pansharpening techniques show results comparable to those of specialized HS pansharpening solutions. It is the case of methods such as MTF-GLP-FS, MTF-GLP-HPM-R, PRACS, GSA, TV. However, it is worth noticing that these methods, originally designed for a smaller number of bands that are well correlated with the PAN, may produce worse results on the bands beyond the PAN spectral bandwidth. For example, MTF-GLP-HPM performs well in the visible range but tends to produce oversharpened results in the NIR and SWIR wavelenghts. Similar observations can be made for methods such as BT-H and BDSD-PC. The take home message is therefore that yes, MS-native solutions can be extended to the HS case taking care of the quality unbalance that may arise in the spectral domain.